OpenStudy (anonymous):

use the quadratic formula to solve each equation. 3x^2+39x+108=0

OpenStudy (anonymous):

x=-9,-4

OpenStudy (anonymous):

When you simply the equation further, you get : \[x ^{2} + 13x + 36 = 0\] Assuming you know the quadratic formula : a = 1 b = 13 c= 36 \[\frac{ -13 \pm \sqrt{169 - 4 \times 1 \times 36} }{ 2\times 1}\] Can you solve it now?

OpenStudy (anonymous):

yeah thank you!